OpenStudy (anonymous):

4/3+2/3q=-5/3-1/3-1/3q-4 solve equation when you are putting in a number to the computer what hey do you use for the negative symbol?

OpenStudy (ash2326):

For negative symbol press the key next to 0 on the keyboard, -

OpenStudy (ash2326):

in the last term q-4 is in the numerator or denominator

OpenStudy (anonymous):

thank you for the info about the key, the q is in the denominator

OpenStudy (anonymous):

did you mean \[\frac{4}{3}+\frac{2}{3q}=-\frac{5}{3}-\frac{1}{3}-\frac{1}{3q}-4\]

OpenStudy (ash2326):

ok 4/3 +2/3q=-5/3-1/3-1/3q-4 2/3q+1/3q=-5/3-1/3-4-4/3 3/3q=-(5/3+1/3+4/3+4) 1/q=-(10/3+4) 1/q=-(22/3) q=-3/22

OpenStudy (anonymous):

yes this is what I mean

OpenStudy (anonymous):

thank you